# Near-Field Localization via Reconfigurable Antennas

### 17 Mar 2026 v1

@@ Abstract
Reconfigurable antennas (RAs) utilize the electromagnetic (EM) domain to provide dynamic control over antenna radiation patterns, which offers an effective way to enhance power efficiency in wireless links. Unlike conventional arrays with fixed element patterns, RAs enable on-demand beam-pattern synthesis by directly controlling each antenna’s EM characteristics. While existing research on RAs has primarily focused on improving spectral efficiency, this paper explores their application for downlink localization. Moreover, the majority of existing works focus on far-field scenarios with little attention on near-field (NF). Motivated by these gaps, we consider a synthesis model in which each antenna generates desired beampatterns from a finite set of EM basis functions. We then formulate a joint optimization problem for the baseband (BB) and EM precoders with the objective of minimizing the user equipment (UE) position error bound (PEB) in NF conditions. Our analytical derivations and extensive simulation results demonstrate that the proposed hybrid precoder design for RAs significantly improves UE positioning accuracy compared to traditional non-reconfigurable arrays.
@@ I Introduction
Localization has become widespread in applications such as wireless networks, augmented reality, and the internet of things (IoT) [4, 8] within the framework of integrated sensing and communication (ISAC). Localization methods are commonly categorized by whether the source lies in the receiver’s far-field (FF) or near-field (NF) region [11]. Sources in the NF region produce spherical wavefronts at the receiver, whereas FF sources are seen as plane waves and are thus described only by angle-of-arrival (AOA) or angle-of-departure (AOD) [5, 16]. Therefore, algorithms developed under the FF assumption are generally unsuitable for NF source localization.
Both FF and NF localization have been extensively investigated across diverse scenarios in the literature (e.g., [11, 15, 21, 10, 8, 5, 7] and the references therein). Most prior works have emphasized beamforming via per-element phase and amplitude control to steer power toward desired directions. A major limitation in this body of work, however, is the common practice of modeling antenna arrays using elements with fixed, frequently identical radiation patterns [14]. The constraints imposed by conventional pattern-fixed antennas are becoming increasingly relevant in emerging applications, such as NF communications [14, 23]. In particular, very narrow element patterns yield high gain only within limited angular sectors and suffer steep loss near endfire, while broad element patterns improve coverage but reduce peak gain.
@@
Recent advances in RA technology have shown growing interest in their integration within antenna arrays [22]. While several works have demonstrated the spectral-efficiency benefits of RA arrays[14, 20], their potential for wireless sensing, particularly for localization, has received limited attention. Furthermore, the majority of prior studies in both communications and sensing assume FF conditions [9]. Motivated by these gaps, this paper targets localization in NF scenarios. In particular, the main contributions of this work are summarized as follows:
@@ II-A System Model
As depicted in Fig. 1, we analyze a multiple input single output (MISO) wireless system where the base station (BS) is equipped with a uniform planar array (UPA). This array consists of $M=M^{h}M^{v}$ reconfigurable antennas (RAs), with $M^{h}$ and $M^{v}$ representing the number of rows and columns, respectively. The position of the center of the BS and the $m$-th RA are denoted by $\mathbf{p}_{b}$ and $\mathbf{p}_{b,m}$, respectively. Each RA is capable of synthesizing a desired pattern. The UE is assumed to have a single isotropic antenna located at the position $\mathbf{p}_{u}$. Assuming each multipath component (MPC) arises from a single reflection at point scatterers, we model multipath using $I$ scatterers with positions $\{\mathbf{p}_{s}^{(i)}\}_{i=1}^{I}$ based on a geometric channel model, which will be detailed in the subsequent subsection. For localization, the BS transmits $N_{t}$ narrowband, single-carrier pilots to the UE.
@@ II-B Signal Model
Under the assumption of a geometric channel model that includes a line-of-sight (LoS) path and $I$ non-line-of-sight (NLoS) paths, the spatial-domain channel vector $\mathbf{h}_{t}$ is given by [20, 22]:

### 22 Mar 2026 v2

@@ Abstract
Reconfigurable antennas (RAs) utilize the electromagnetic (EM) domain to provide dynamic control over antenna radiation patterns, which offers an effective way to enhance power efficiency in wireless links. Unlike conventional arrays with fixed element patterns, RAs enable on-demand beam-pattern synthesis by directly controlling each antenna’s EM characteristics. While existing research on RAs has primarily focused on improving spectral efficiency, this paper explores their application for downlink localization. Moreover, the majority of existing works focus on far-field scenarios with little attention on near-field (NF). Motivated by these gaps, we consider a synthesis model in which each antenna generates desired beampatterns from a finite set of EM basis functions. We then formulate a joint optimization problem for the baseband ($B$) and EM precoders with the objective of minimizing the user equipment (UE) position error bound (PEB) in NF conditions. Our analytical derivations and extensive simulation results demonstrate that the proposed hybrid precoder design for RAs significantly improves UE positioning accuracy compared to traditional non-reconfigurable arrays.
@@ I Introduction
Localization has become widespread in applications such as wireless networks, augmented reality, and the internet of things (IoT) [4, 8] within the framework of integrated sensing and communication (ISAC). Localization methods are commonly categorized by whether the source lies in the receiver’s far-field ($F$) or near-field (NF) region [11]. Sources in the NF region produce spherical wavefronts at the receiver, whereas $F$ sources are seen as plane waves and are thus described only by angle-of-arrival (AOA) or angle-of-departure (AOD) [5, 16]. Therefore, algorithms developed under the $F$ assumption are generally unsuitable for NF source localization.
Both $F$ and NF localization have been extensively investigated across diverse scenarios in the literature (e.g., [11, 15, 21, 10, 8, 5, 7] and the references therein). Most prior works have emphasized beamforming via per-element phase and amplitude control to steer power toward desired directions. A major limitation in this body of work, however, is the common practice of modeling antenna arrays using elements with fixed, frequently identical radiation patterns [14]. The constraints imposed by conventional pattern-fixed antennas are becoming increasingly relevant in emerging applications, such as NF communications [14, 23]. In particular, very narrow element patterns yield high gain only within limited angular sectors and suffer steep loss near endfire, while broad element patterns improve coverage but reduce peak gain.
@@
Recent advances in RA technology have shown growing interest in their integration within antenna arrays [22]. While several works have demonstrated the spectral-efficiency benefits of RA arrays[14, 20], their potential for wireless sensing, particularly for localization, has received limited attention. Furthermore, the majority of prior studies in both communications and sensing assume $F$ conditions [9]. Motivated by these gaps, this paper targets localization in NF scenarios. In particular, the main contributions of this work are summarized as follows:
@@ II-A System Model
As depicted in Fig. 1, we analyze a multiple input single output (MISO) wireless system where the base station (BS) is equipped with a uniform planar array (UPA). This array consists of $M=M^{h}M^{v}$ reconfigurable antennas (RAs), with $M^{h}$ and $M^{v}$ representing the number of rows and columns, respectively. The position of the center of the BS and the $m$-th RA are denoted by $\mathbf{p}_{b}$ and $\mathbf{p}_{b,m}$, respectively. Each RA is capable of synthesizing a desired pattern. The UE is assumed to have a single isotropic antenna located at the position $\mathbf{p}_{u}$. Assuming each multipath component (MPC) arises from a single reflection at point scatterers, we model multipath using II scatterers with positions $\{\mathbf{p}_{s}^{(i)}\}_{i=1}^{I}$ based on a geometric channel model, which will be detailed in the subsequent subsection. For localization, the BS transmits $N_{t}$ narrowband, single-carrier pilots to the UE.
@@ II-B Signal Model
Under the assumption of a geometric channel model that includes a line-of-sight (LoS) path and II non-line-of-sight (NLoS) paths, the spatial-domain channel vector $\mathbf{h}_{t}$ is given by [20, 22]: